OpenStudy (anonymous):

HeLP ? 3 = –(–y + 6)

OpenStudy (caozeyuan):

solve the equation?

OpenStudy (anonymous):

yes

OpenStudy (caozeyuan):

so you need o expand the bracket

OpenStudy (anonymous):

A. –9 B. –3 C. 3 D. 9

OpenStudy (caozeyuan):

so negative of negative y is positve y and negative of positive 6 is negative 6

OpenStudy (caozeyuan):

so 3=y-6

OpenStudy (caozeyuan):

then move 6 to the other side

OpenStudy (caozeyuan):

now y=3+6=9

OpenStudy (anonymous):

oh so its 9 the answer

OpenStudy (caozeyuan):

thats rihgt

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

one more ?

OpenStudy (caozeyuan):

sure

OpenStudy (anonymous):

(x-16)+5=-15

OpenStudy (anonymous):

@caozeyuan

OpenStudy (caozeyuan):

can you try it by yourself first?

OpenStudy (anonymous):

k

OpenStudy (caozeyuan):

do what I ddi in the last example

OpenStudy (anonymous):

-4 ?

OpenStudy (anonymous):

@caozeyuan

OpenStudy (caozeyuan):

looks like you mastered the step for solving basic algbric eqautions!

OpenStudy (anonymous):

@caozeyuan taught me :D !!!!!!

OpenStudy (caozeyuan):

I am happy to teach anyone math, or other sciences

OpenStudy (caozeyuan):

tag em if you have problems in the future